A spacecraft passes a spherical space station. Observers in the spacecraft measure the station’s minor axis as 145 km and the major axis as 190 km.

(a) How fast is the spacecraft travelling relative to the space station?

(b) Why does the station not look like a sphere to the observers in the spacecraft?


1
Expert's answer
2021-06-21T16:29:38-0400

Gives

Major x-axis (a)=190km

Minor x-axis (b)=145km

Part (a)


"V_{minor}=\sqrt{gR}=\sqrt{gb}=\sqrt{9.8\times145}=37.70km/sec"

"V_{major}=\sqrt{gR}=\sqrt{ga}=\sqrt{9.8\times 190}=43.15km/sec"

Space craft move in minor x-axis velocity Vminor=37.70km/sec

Space craft move in major x-axis velocity

Vmajor=43.15km/sec


"V_{minor}<V_{major}"

Major x-axis velocity is greater than minor x-axis Velocity



"V_{rel}=V_{major}-V_{minor}=43.15-37.70 =5.45km\/sec"

station not look like a sphere to the observers in the spacecraft

Space craft move very long distance of eart orbital

Then not look like a sphere to the observers